# Community-level Contagion among Diverse Financial Assets

### v1

@@ 6 Conclusion
In this study, we discover contagion effects in a broad financial market comprising of stocks, cryptocurrencies and US ETFs during a period when a number of impactful events were taking place. Our work contributes to the existing literature in different ways. First and foremost, the contagion is examined at the community (e.g. a group of assets) level rather than focusing solely on individual asset relationships. Next, we segment the initial full period in question into distinct sub-periods based on major financial events, allowing for a detailed exploration of contagion dynamics under varying market conditions. Moreover, unlike many prior studies that concentrate on a single type of investment vehicle, our analysis spans multiple types, offering a more comprehensive and cross-sectional view of financial contagion.
Observing intra- and inter-community contagion under various market conditions, including the turbulent time caused by the Covid-19 pandemic, bullish phases and periods of political conflict, we find that contagion density both within a community and between two communities tends to intensify during times of market uncertainty, as might be expected. This holds true regardless of whether the uncertainty stems from negative events such as the pandemic or positive developments such as strong market rallies. Notably, the degree of density appears to be proportional with the severity of market uncertainty. In contrast, the magnitude of contagion remains relatively stable across all market conditions. From a systemic risk perspective, this result indicates that market uncertainty is characterized more by the widespread diffusion of shocks than by the intensity of any single spillover. Therefore, ones should pay attention to the expansion of contagion pathways, rather than focusing solely on impact magnitudes.
@@
For the second research question, our analysis of directional contagion density across community pairs reveal that certain communities, particularly those composed of assets from the Information Technology sector, act as dominant contagion transmitters in the network, during each sub-period. This strong influence is likely driven by the sector’s strong growth prospects and substantial market capitalization. Additionally, communities forming mostly from Healthcare or Financials assets also emerge as dominant transmitters during specific sub-periods, possibly caused by prevailing market conditions. This result emphasizes the necessity for investors to monitor the Information Technology sector closely as its price movements or shocks may ripple through a wide range of assets. Furthermore, major events and their associated business sectors should be carefully monitored since they pose a heightened risk of widespread contagion throughout the market coming from the sectors.
While this study provides valuable insights into the financial contagion, several rooms for future research remain open. Firstly, the analysis could be extended to a broader market by including a larger set of assets and more types of investment vehicle. Secondly, expanding the analysis to include longer time horizons or more granular time windows could offer a better understanding of how contagion evolves through different market regimes. Thirdly, incorporating exogenous variables such as macroeconomic indicators, monetary policy changes and sentiment scores from news and social media may help uncover the drivers behind contagion effects. Fourthly, another promising direction is to develop forecasting frameworks that incorporate contagion information to enhance forecasting accuracy.
Acknowledge
This research was conducted with the financial support of Research Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT Centre at Dublin City University. ADAPT, the Research Ireland Centre for AI-Driven Digital Content Technology, is funded through the Research Ireland Centres Programme. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 7 Conclusion
In this study, we discover contagion effects in a broad financial market comprising stocks, cryptocurrencies and US ETFs during a period when a number of impactful events were taking place. Our work contributes to the existing literature in different ways. First and foremost, the contagion is examined at the community (e.g. a group of assets) level rather than focusing solely on individual asset relationships. Next, we segment the initial full period in question into distinct sub-periods based on major financial events, allowing for a detailed exploration of contagion dynamics under varying market conditions. Moreover, unlike many prior studies that concentrate on a single type of investment vehicle, our analysis spans multiple types, offering a more comprehensive and cross-sectional view of financial contagion.
Observing intra- and inter-community contagion under various market conditions, including the turbulent time caused by the Covid-19 pandemic, bullish phases and periods of political conflict, we find that contagion density both within a community and between two communities tends to intensify during times of market uncertainty, as might be expected. This holds true regardless of whether the uncertainty stems from negative events such as the pandemic or positive developments such as strong market rallies. Notably, the degree of density appears to be proportional to the severity of market uncertainty. In contrast, the magnitude of contagion remains relatively stable across all market conditions. From a systemic risk perspective, this result indicates that market uncertainty is characterized more by the widespread diffusion of shocks than by the intensity of any single spillover. Therefore, one should pay attention to the expansion of contagion pathways, rather than focusing solely on impact magnitudes.
@@
For the second research question, our analysis of directional contagion density across community pairs reveals that certain communities, particularly those composed of stocks and US ETFs dominated by holdings in the Information Technology sector, act as dominant contagion transmitters in the network, during each sub-period. This strong influence is likely driven by the sector’s strong growth prospects and substantial market capitalization. Additionally, communities forming mostly from Healthcare, Financials, Energy and cryptocurrency assets also emerge as dominant transmitters during specific sub-periods, possibly caused by prevailing market conditions. This result emphasizes the necessity for investors to monitor the Information Technology sector closely as its price movements or shocks may ripple through a wide range of assets. Furthermore, major events and their associated business sectors should be carefully monitored since they pose a heightened risk of widespread contagion throughout the market coming from the sectors.
While this study provides valuable insights into the financial contagion, several rooms for future research remain open. Firstly, the analysis could be extended to a broader market by including a larger set of assets and more types of investment vehicle. Secondly, expanding the analysis to include up-to-date data could provide additional insights to this study. Thirdly, incorporating exogenous variables such as macroeconomic indicators, monetary policy changes and sentiment scores from news and social media may help uncover the drivers behind contagion effects. Fourthly, another promising direction is to develop forecasting frameworks that incorporate contagion information to enhance forecasting accuracy.
CRediT authorship contribution statement
An Pham Ngoc Nguyen: Writing – review & editing, Writing – original draft, Visualization, Validation, Software, Project administration, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Marija Bezbradica: Writing – review & editing, Validation, Supervision, Resources, Project administration, Investigation, Funding acquisition, Conceptualization. Martin Crane: Writing – review & editing, Validation, Supervision, Resources, Project administration, Investigation, Funding acquisition, Conceptualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This research was conducted with the financial support of Taighde Éireann-Research Ireland under Grant Agreement No. 13/RC/2106_P2 at the ADAPT Centre at Dublin City University. ADAPT, the Research Ireland Centre for AI-Driven Digital Content Technology, is funded through the Research Ireland Centres Programme. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Appendix A. Supplementary data
Supplementary material related to this article can be found online at
https://doi.org/10.1016/j.chaos.2025.117858.
Data availability
Data is available in my Github repo:
https://github.com/NguyenPhamNgocAn/Community-Level-Contagion-in-Financial-Markets.